OpenStudy (anonymous):

The sum of three consecutive integers is 267. What is the largest integer?

OpenStudy (raden):

can you solve for x, first ? x + x + 1 + x + 2 = 267 after you get x, the answer is x + 2 = ... ?

OpenStudy (anonymous):

Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now

OpenStudy (anonymous):

Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now

OpenStudy (anonymous):

Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now

OpenStudy (raden):

J = 10 + T ... (1) J - 3 = 2(T - 3) = 2T - 6 ... (2) sutract the 2nd equation by the 1st, solve for T (Tim's old)